OpenStudy (rainbow_rocks03):

What is the product of 7/9 and 4/5?

OpenStudy (alexandervonhumboldt2):

\[\frac{ 7 }{ 9}*\frac{ 4 }{ 5 }=\frac{ 7*4 }{ 9*5 }=simpify\]